OpenStudy (anonymous):

Calc problem: Help please A cable hangs between two poles of equal height and 33 feet apart. At a point on the ground directly under the cable and x feet from the point on the ground halfway between the poles the height of the cable in feet is h(x)=10 +(0.3)( x^{1.5}) . The cable weighs 12.3 pounds per linear foot. Find the weight of the cable.

OpenStudy (anonymous):

If you put the function h(x) into the arc length formula you get its length, then multiply by 12.3 to get the total weight.

OpenStudy (anonymous):

\[L= \int\limits_{a}^{b} \sqrt{1+(10+(0.3)(x ^{1.5})} *(12.3)\]

OpenStudy (anonymous):

what would the value of x be?

OpenStudy (anonymous):

under the radical it should be 1+(h'(x))^2 not 1+h(x) so find h'(x) and square it. integrate with respect to x and use a=0 and b=33 feet. Then multiply by 12.3

OpenStudy (anonymous):

\[L= \int\limits_{0}^{33} \sqrt{1+(\frac{ 9\sqrt{x} }{ 20 }})^2\] *then multiply by 12.3 i got 821.763 @slickwilly

OpenStudy (anonymous):

I'm not sure about that integral. These ones are usually hard to integrate. Maybe someone else will figure it out but I can't and I'm pretty sure 821 ft is way too big.

OpenStudy (anonymous):

help? @campbell_st (:

OpenStudy (campbell_st):

sorry can't offer anything

OpenStudy (anonymous):

Thanks for looking into the problem though!

OpenStudy (anonymous):

I still got a very large answer... @campbell_st

OpenStudy (anonymous):

I just noticed h(x) goes from the midpoint to one of the ends so I think you need to integrate from 0 to 33/2 and then double the result.

OpenStudy (anonymous):

I still got a large answer in the thousands.

OpenStudy (campbell_st):

ok... just had a quick read you need \[h'(x) = 0.45x^{0.5}\] then the parabolic arc length is \[L = \int\limits_{a}^{b} \sqrt{1 + (h'(x))^2} dx\] which means you are looking at \[L = \int\limits_{0}^{33} \sqrt{1 + 0.45^2x} dx... or L = \int\limits_{0}^{33}\sqrt{1 + 0.2025x} dx\]

OpenStudy (anonymous):

would I just evaluate both integrals and guess/check?

OpenStudy (anonymous):

do I have to multiply L by 12.3

OpenStudy (anonymous):

The last integral looks right but do it from 0 to 33/2 and then double it. Then multiply by 12.3

OpenStudy (anonymous):

okay I got 19.04 doubling it I got 38.08 then multiplying it by 12.3 I got 468.384 and that was wrong... ;/ @slickwilly

OpenStudy (anonymous):

I think I'm out of ideas for this one. Good luck.